A guitar string with a linear density of 2.0 g/m is stretched between supports that are 62 cm apart. The string is observed to form a standing wave with three antinodes when driven at a frequency of 440 Hz.

What is the frequency of a standing wave with five antinodes on this string? What is the tension of the string?
Answer #1

A)

For 3 antinodes,

3f = 440

Frequency, f = 440/3 = 146.67 Hz

Frequency of frequency with 5 antinodes, f' = 5 x 146.67 = 733.33 Hz

B)

Tension, T = f^2 x 4L^2 x u

T = 146.67^2 x 4 x 0.62^2 x 0.002 = 66.15 N

Comment in case any doubt please rate my answer.....